### v2

# Superamalgamation for modal lattices via non-distributive dualities

Abstract
We show that the variety of modal lattices has the superamalgamation property. As a consequence, we obtain that the weak positive modal logic has the Craig interpolation property. Our proof employs the recent duality for modal lattices based on modal L-spaces. Moreover, we extend this result to a number of other weak positive modal logics axiomatized by modal axioms corresponding to universal Horn sentences.
1 Introduction
Superamalgamation is an important property in universal algebra. It is especially crucial in algebraic logic, where it provides an algebraic counterpart of the notion of Craig interpolation. More formally, the standard connection is as follows: a logic $\mathcal{L}$ has the Craig interpolation property if and only if its corresponding variety of algebras $\mathsf{K}(\mathcal{L})$ has the superamalgamation property. We refer to the recent handbook on interpolation [23] for the current state of the art on Craig interpolation, as well as to Chapters [2, 10, 21] of this handbook for detailed discussions of the connections between (super)amalgamation and Craig interpolation.
Maksimova’s classical result [18] states that there are exactly seven consistent superintuitionistic logics with Craig interpolation. This result was proved by characterizing all seven nontrivial varieties of Heyting algebras with the amalgamation property (in varieties of Heyting algebras, the notions of amalgamation and superamalgamation coincide); for the proof we refer, for example, to [5, Section 14.4]. Using similar techniques, Maksimova [19] also showed that there are at most 37 modal logics extending $\mathsf{S4}$ that have the Craig interpolation property (see, e.g., [5, 10, 11]).
In this paper, we study a natural version of the Craig interpolation property for modal logics lacking an implication connective which are moreover not necessarily distributive. Such logics were named weak positive modal logics in [1], where “weak” refers to the possible lack of the distributivity law. Most such results are proved by establishing that the class of modal lattices has superamalgamation. In this context, Jónsson [17] (see also [16]) proved that the variety of lattices has the amalgamation property. Subsequently, [7] showed that the only nontrivial varieties of lattices with the amalgamation property are the variety of all lattices and the variety of distributive lattices. Note also that the variety of distributive lattices lacks the superamalgamation property, making the possible lack of distributivity a natural assumption.
Jónnson’s proof is algebraic, making use of filters on lattices; whilst this allows for ample generalization, when considering modalities, these might not be definable on the filters. To overcome this limitation, we approach this problem through duality theory. Duality is a powerful tool that provides topological representations of algebraic structures. For dualities for Boolean algebras, distributive lattices, Heyting algebras, and modal algebras we refer to [6, 9, 12, 24]. In all these cases, duality yields a transparent dual proof of the amalgamation property: briefly, one dualizes all objects and morphisms and then constructs an amalgam using the dual construction of a pullback (in an appropriate category). For proofs in the case of modal algebras, we refer to [2, 24].
In the setting of modal lattices that are not necessarily distributive, a new duality, resembling the classical dualities in the distributive setting, was recently developed in [1]. We briefly recall the main points of this duality. Given a lattice $L$, one considers the set of all filters of $L$, equipped with the patch topology of the standard Stone topology, ordered by inclusion of filters. This order is easily seen to form a meet semilattice, the topology is compact, and whenever $x\not\preceq y$, then $x$ and $y$ are separated by a clopen filter — this yields the structure of an L-space. Given an L-space $(X,\tau)$, one obtains a lattice by taking the lattice of all clopen filters of $X$. In this way, every lattice is represented as the lattice of clopen filters of its dual L-space. This correspondence can be lifted to a representation and duality of modal lattices [1].
Using this duality, and essentially the same ideas as those employed in the distributive setting, we show that the variety of modal lattices has the superamalgamation property. Consequently, the corresponding weak positive modal logic has the Craig interpolation property. There are both similarities and differences of this approach with the classical case. In the classical setting, one works with the complex algebra of a frame (i.e., the powerset algebra or the algebra of all upsets in the distributive and Heyting cases); these correspond to canonical extensions of modal algebras, distributive lattices or Heyting algebras [9, 12, 24]. The analogue in our setting is the modal lattice of all filters of a meet semilattice, which in turn corresponds to the $\Pi_{1}$-completions of modal lattices [1].
Moreover, as in the classical case, our proof extends to all logics whose frame correspondents are axiomatized by universal Horn sentences [2, 24]. We give several illustrative examples in Section 4. Surprisingly, the weak positive modal logic axiomatized by the Church–Rosser $\mathsf{.2}$ axiom $\Diamond\Box p\trianglelefteq\Box\Diamond p$ also has the Craig interpolation property (see Theorem 4.19). This should be contrasted with the classical case, where the modal logic $\mathsf{K.2}$ lacks the Craig interpolation property [20, Section 5.6.2]. Moreover, we show that despite lacking superamalgamation, the logic of distributive lattices enjoys the Craig interpolation property.
The paper is organized as follows. In Section 2, we first recall the syntax and algebraic semantics of weak positive modal logic and then introduce the algebraic counterparts of interpolation related properties. In Section 3, we present a new proof of the superamalgamation property for lattices, first recalling the relevant duality and then using it to establish our result. In Section 4, we build on the results of Section 3 by adding modalities. This is followed by a discussion of the Craig interpolation property for certain extensions of weak positive modal logic. In the final section, we offer some concluding remarks.
2 Preliminaries
2.1 Modal lattices and weak positive modal logic
We assume the reader is familiar with basic concepts from lattice theory [13] and universal algebra [3]. The language of modal lattices is given by the BNF:
$p\mid\top\mid\bot\mid\phi\land\psi\mid\phi\lor\psi\mid\Box\phi\mid\Diamond\phi.$
The language of lattices is obtained by omitting all formulas that contain modalities $\Box$ or $\Diamond$. We assume all lattices to be bounded.
Definition 2.1.
Given a set of propositional letters $P$, we let $Fm_{P}$ be the set of formulas generated from those propositional letters using the language of modal lattices.
A consequence pair is an expression of the form $\phi\trianglelefteq\psi$, with $\phi,\psi\in Fm_{P}$.
Definition 2.2.
The deduction rules and axioms for consequence pairs are as follows:
$\trianglelefteq\top, p\bot\trianglelefteq p$, top and bottom
[…]
$p\land q\trianglelefteq p$, $p\land q\trianglelefteq q$, left conjunction rules
$p\trianglelefteq p\lor q$, $q\trianglelefteq p\lor q$, right disjunction rules
[…]
$\top\trianglelefteq\Box\top$, $\top\trianglelefteq\Diamond\top$, modal top
[…]
$\Box p\land\Box q\trianglelefteq\Box(p\land q)$, $\Diamond p\land\Box q\trianglelefteq\Diamond(p\land q)$, linearity and duality
We denote by $\mathcal{L}$ the smallest set of consequence pairs closed under uniform substitution and the axioms and rules listed above.
If $\Gamma$ is a set of consequence pairs, $\mathcal{L}(\Gamma)$ denotes the smallest set of modal consequence pairs containing $\Gamma$ and closed under the axioms and rules listed above. Such a set is called a weak positive modal logic, or a logic for short. We write $\phi\vdash_{\Gamma}\psi$ if $\phi\trianglelefteq\psi\in\mathcal{L}(\Gamma)$.
Definition 2.3.
A modal lattice is a tuple $(A,\top,\bot,\land,\lor,\Box,\Diamond)$ such that $(A,\top,\bot,\land,\lor)$ is a bounded lattice, that the following identities hold:
$\top$ $\approx\Box\top,$ $\top$ $\approx\Diamond\top,$
$\Box a\land\Box b$ $\approx\Box(a\land b),\qquad$ $\qquad\Diamond a\lor\Diamond b$ $\leq\Diamond(a\lor b),$
$\Diamond a\land\Box b$ $\leq\Diamond(a\land b).$
We denote the category of modal lattices by $\mathsf{MLat}$, and the category of lattices by $\mathsf{Lat}$.
Remark 2.4.
We note that the dual Dunn axiom, $\Box(a\lor b)\leq\Box a\lor\Diamond b$ need not hold [1, Remark 4.8], and similarly the symmetric normality axiom for $\Diamond$, namely $\Diamond(a\lor b)\leq\Diamond a\lor\Diamond b$, need not hold [1, Example 4.41].
We can use modal lattices as models of weak positive modal logic: the set $Fm_{P}$ comes naturally equipped with a modal lattice structure, and so given a modal lattice $A$, we call a homomorphism $\sigma\colon Fm_{P}\to A$ a valuation.
Definition 2.5.
We say that a modal lattice $A$ validates a consequence pair $\phi\trianglelefteq\psi$, written $A\vDash\phi\trianglelefteq\psi$, if for every valuation $\sigma\colon Fm_{P}\to A$, we have $\sigma(\phi)\leq\sigma(\psi)$.
A variety $\mathsf{K}$ validates a consequence pair $\phi\trianglelefteq\psi$, written $\vDash_{\mathsf{K}}\phi\trianglelefteq\psi$, if for all $A\in\mathsf{K}$, we have $A\vDash\phi\trianglelefteq\psi$.
Definition 2.6.
Given a set of consequence pairs $\Gamma$, we denote by $\mathsf{K}(\Gamma)$ the variety of lattices that validate all the pairs in $\Gamma$. We use $\vDash_{\Gamma}$ as a shorthand for $\vDash_{\mathsf{K}(\Gamma)}$.
Given a variety, we denote by $\mathcal{L}(\mathsf{K})$ the set of consequence pairs that are validated by $\mathsf{K}$. This gives a correspondence between varieties of modal lattices and modal logics.
Definition 2.7.
Given a variety $\mathsf{K}$, we denote by $F_{\mathsf{K}}(P)$ the free algebra on the set of generators $P$ in $\mathsf{K}$. When the variety is of the form $\mathsf{K}(\Gamma)$, we denote the free algebra by $F_{\Gamma}(P)$. It is obtained from $Fm_{P}$ by identifying pairs $\phi,\psi$ such that $\phi\vdash_{\Gamma}\psi$ and $\psi\vdash\phi$.
Remark 2.8.
For all of the objects defined above, we drop the subscript $\Gamma$ when $\Gamma$ is the empty set.
Free algebras play a key role in the algebraic completeness theorems, and they will be used heavily throughout our work. We collect some basic facts concerning them.
Proposition 2.9.
  1. (i)
    Given $\phi\in Fm_{P}$, and $\mathsf{K}$ a variety of modal lattices, we can identify $\phi$ with $\phi]_{\mathsf{K}}$, its equivalence class in the free algebra $F_{\mathsf{K}}(P)$. Under this identification, we have that $\phi\vdash_{\mathsf{K}}\psi$ if and only if $\phi\leq\psi$ holds in $F_{\mathsf{K}}(P)$.
  2. (ii)
    If $\sigma_{1}\colon F_{\mathsf{K}}(\overline{p},\overline{q})\to A$ and $\sigma_{2}\colon F_{\mathsf{K}}(\overline{p},\overline{r})\to A$ are homomorphisms which agree on $\overline{p}$, then there is a unique morphism $\sigma\colon F_{\mathsf{K}}(\overline{p},\overline{q},\overline{r})\to A$ which agrees with $\sigma_{1}$ and $\sigma_{2}$ on the restrictions.
The following theorem is proved in [1, Theorem 4.13].
Theorem 2.10 (Algebraic Completeness Theorem).
For $\phi,\psi\in Fm_{P}$ and $\Gamma$ a set of formulas, we have $\phi\vdash_{\Gamma}\psi$ iff $\vDash_{\Gamma}\phi\trianglelefteq\psi$.
2.2 Craig interpolation and Superamalgamation
We move on to studying Craig interpolation. Usually Craig interpolation is defined for languages containing an implication connective. The following definition is similar to this by treating the $\vdash$ as a form of implication:
Definition 2.11.
We say that a logic $\mathcal{L}(\Gamma)$ has the Craig interpolation property if for all $\phi\in Fm_{\overline{p},\overline{q}}$ and $\psi\in Fm_{\overline{p},\overline{r}}$ such that $\phi\vdash_{\Gamma}\psi$, there is a formula $\chi\in Fm_{\overline{p}}$ such that $\phi\vdash_{\Gamma}\chi$ and $\chi\vdash_{\Gamma}\psi$. Such a formula $\chi$ is called an interpolant.
Definition 2.12.
Let $\mathsf{K}$ be a variety of modal lattices. A V-formation is a tuple of lattices $K,L_{1},L_{2}\in\mathsf{K}$ and embeddings $h_{i}\colon K\to L_{i}$, as in Figure 1(a)
A variety $\mathsf{K}$ of modal lattices has the amalgamation property if for every V-formation, there is a lattice $M$, and injective homomorphisms $p_{i}\colon L_{i}\to M$ such that $p_{1}\circ h_{1}=p_{2}\circ h_{2}$, or in other words, such that the diagram in Figure 1(b) commutes. The lattice $M$ is called an amalgam.
[…]
(a) V-formation
[…]
(b) Amalgam
Figure 1: Amalgamation of lattices
A variety $\mathsf{K}$ of modal lattices has the superamalgamation property if for every V-formation, there is an amalgam such that $p_{1}(a)\leq p_{2}(b)$ for $a\in L_{1}$ and $b\in L_{2}$ implies the existence of $c\in K$ such that $a\leq h_{1}(c)$ and $h_{2}(c)\leq b$. The lattice $M$ is called a superamalgam.
Definition 2.13.
A variety $\mathsf{K}$ of modal lattices has epimorphism surjectivity if every epimorphism in the category corresponding to the variety is onto.
Remark 2.14.
The epimorphism surjectivity property is the algebraic equivalent of the Beth definability property. However, we decide to not go into the details of the Beth definability property here (see [15] for details).
Lemma 2.15.
Let $\mathsf{K}$ be a variety of lattices. The following are equivalent:
  1. (i)
    the V-formation in Figure 2(a) superamalgamates;
  2. (ii)
    the amalgamation diagram in Figure 2(b) is a superamalgamation;
  3. (iii)
    the logic $\mathcal{L}(\mathsf{K})$ has the Craig interpolation property for $\phi\in Fm_{\overline{p},\overline{q}}$ and $\psi\in Fm_{\overline{p},\overline{r}}$. That is, any pair $\phi\in Fm_{\overline{p},\overline{q}}$ and $\psi\in Fm_{\overline{p},\overline{r}}$ such that $\vDash_{\mathsf{K}}\phi\trianglelefteq\psi$ has an interpolant.
$F_{\mathsf{K}}(\overline{p},\overline{q})$$F_{\mathsf{K}}(\overline{p},\overline{r})$$F_{\mathsf{K}}(\overline{p})$}
(a) V-formation of free algebras
$F_{\mathsf{K}}(\overline{p},\overline{q},\overline{r})$$F_{\mathsf{K}}(\overline{p},\overline{q})$$F_{\mathsf{K}}(\overline{p},\overline{r})$$F_{\mathsf{K}}(\overline{p})$}
(b) Amalgam of free algebras
Figure 2: Craig interpolation and superamalgamation on free algebras
Proof.
That (iii) implies (ii) follows simply by unfolding the definitions. That (ii) implies (i) is trivial. Let us prove that (i) implies (iii). Assume that we have a superamalgam $M$ with two monos $\sigma_{1}\colon F_{\mathsf{K}}(\overline{p},\overline{q})\to M$ and $\sigma_{2}\colon F_{\mathsf{K}}(\overline{p},\overline{r})\to M$ which agree on $F_{\mathsf{K}}(\overline{p})$. Assume that $\phi\in F_{\mathsf{K}}(\overline{p},\overline{q})$ and $\psi\in F_{\mathsf{K}}(\overline{p},\overline{r})$ are such that $\phi\vdash_{\mathsf{K}}\psi$. Then by Proposition (ii), there is a map $\sigma\colon F_{\mathsf{K}}(\overline{p},\overline{q},\overline{r})\to M$ extending $\sigma_{1},\sigma_{2}$. As $\phi\vdash_{\mathsf{K}}\psi$, we have $\sigma_{1}(\phi)=\sigma(\phi)\leq\sigma(\psi)=\sigma_{2}(\psi)$. Therefore by superamalgamation, there is $\chi\in F_{\mathsf{K}}(\overline{p})$ such that $\phi\leq\chi$ in $F_{\mathsf{K}}(\overline{p},\overline{q})$, and $\chi\leq\psi$ in $F_{\mathsf{K}}(\overline{p},\overline{r})$. By Proposition (i), this implies that $\phi\vdash_{\mathsf{K}}\chi$ and $\chi\vdash_{\mathsf{K}}\psi$, thus $\chi$ is the interpolant for $\phi$ and $\psi$. ∎
Recall that a bounded lattice $L$ is called a Heyting algebra if for all $a,b\in L$ there is an element $a\rightarrow b$ such that for all $c\in L$:
$a\land c\leq b\qquad\text{iff}\qquad c\leq a\rightarrow b.$
For basic facts about Heyting algebras see [5, Chapter 7].
Recall that positive logic is the logic of distributive lattices, see, e.g., [4] for more details.
Proposition 2.16.
Positive logic has Craig interpolation.
Proof.
We denote the variety of distributive lattices by $\mathsf{DLat}$. By Lemma 2.15, it is sufficient to show that the class of free distributive lattices has superamalgamation. However, any free distributive lattice is a Heyting algebra, and the inclusion of $\iota\colon F_{\mathsf{DLat}}(\{p_{1},\dots,p_{n}\})\to F_{\mathsf{DLat}}(\{p_{1},\dots,p_{n},p_{n+1}\})$ is a Heyting algebra homomorphism  [^1] For a proof and references to this result, see [22, Prop. 2.12]., and the variety of Heyting algebras has superamalgamation [18]. ∎
The following gives us a natural connection between superamalgamation, Craig interpolation and epimorphism surjectivity:
Theorem 2.17.
Let $\mathsf{K}$ be a variety of modal lattices and $\mathcal{L}(\mathsf{K})$ be the corresponding logic.
  1. (i)
    If $\mathsf{K}$ has superamalgamation, then $\mathcal{L}(\mathsf{K})$ has Craig interpolation.
  2. (ii)
    If $\mathsf{K}$ has superamalgamation, then it has epimorphism surjectivity.
Proof.
(i) This follows from Lemma 2.15 that $\mathcal{L}(\mathsf{K})$ has Craig interpolation.
(ii) This is given in [15, Corollary 2.5.23]. ∎
Remark 2.18.
The converse to (i) is true when $\mathsf{K}$ is a variety of Heyting algebras. Note that this is necessary for Craig interpolation to imply superamalgamation. Indeed, Corollary 2.16 tells us that the logic of distributive lattices has Craig interpolation. However, distributive lattices do not have epimorphism surjectivity, and therefore, fail to have superamalgamation. Indeed, the inclusion from the $3$-chain to the $4$ elements Boolean algebra (see Figure 3) depicted below is an epimorphism, but it is not surjective.
$\bullet\bullet\bullet\bullet\bullet\bullet\bullet$
Figure 3: An epimorphism which is not surjective
3 Superamalgamation of lattices
In this section, we focus on (non modal) lattices. While none of the results presented in this section are novel, their proofs are new, and they serve as a convenient stepping stone for the results in the next section.
3.1 Duality for lattices
We now recall the duality for lattices and modal lattices developed in [1]. As it is obtained by restricting HMS (Hoffman-Mislove-Stralka) duality [14], let us first recall the concepts at play there.
Definition 3.1.
An L-frame is a meet semilattice $(X,1,\curlywedge)$. We denote its underlying order by $\preceq$. Given an L-frame $X$, we denote by $\mathsf{Fil_{F}}(X)$ its set of filters, which is a complete lattice.
Definition 3.2.
An HMS-space is a tuple $(X,1,\curlywedge,\tau)$ such that $(X,1,\curlywedge)$ is an L-frame and $(X,\tau)$ is a compact topological space such that whenever $x\not\preceq y$, there is some clopen filter $U\subseteq X$ such that $x\in U$ and $y\not\in U$. An HMS-morphism is a continuous meet semilattice homomorphism. We write $\mathsf{HMS}$ for the category of HMS-spaces and HMS-morphisms.
Theorem 3.3.
The category $\mathsf{HMS}$ is dually equivalent to the category $\mathsf{MSL}$ of meet semilattices.
Proof.
For a complete proof see [14]. We briefly recall the functors involved in this duality. One part is the assignment $\mathsf{Fil}\colon\mathsf{MSL}\to\mathsf{HMS}$, that sends a lattice to its set of filters. Given a meet semilattice $L$, the set $\mathsf{Fil}(L)$ is a meet semilattice, with meet given by intersection. We equip it with a Stone like topology $\tau$ given by declaring the sets
$\varphi(a)\coloneq\{F\in\mathsf{Fil}(L)\mid a\in F\},$
and its complements to be clopen. We obtain that $(\mathsf{Fil}(L),L,\subseteq,\tau)$ is an HMS-space. Conversely, if $(X,\tau)$ is an HMS-space, the set $\mathsf{ClopFil}(X)$ of clopen filters forms a meet semilattice. Given a meet semilattice homomorphism $h\colon K\to L$, the map $h^{-1}\colon\mathsf{Fil}(L)\to\mathsf{Fil}(K)$ will be a continuous meet semilattice map. And given an HMS-morphism $f\colon X\to Y$, the map $f^{-1}\colon\mathsf{ClopFil}(Y)\to\mathsf{ClopFil}(X)$ will be a meet semilattice homomorphism. ∎
In the setting of weak positive logic, the morphisms of interest between L-frames satisfy more requirements:
Definition 3.4.
An L-morphism between L-frames $X$ and $Y$ is a semilattice homomorphism that satisfies for all $x\in X$ and $y^{\prime},z^{\prime}\in Y$:
  1. (i)
    if $f(x)=1$, then $x=1$;
  2. (ii)
    if $y^{\prime}\curlywedge z^{\prime}\preceq f(x)$, then there are $y,z\in X$ such that $y^{\prime}\preceq f(y)$, $z^{\prime}\preceq f(z)$ and $y\curlywedge z\preceq x$ (see Figure 4).
We denote the category of L-frames and L-morphisms by $\mathsf{LFrm}$.
[…]
Figure 4: The second L-morphism condition
The definition of L-morphism ensures that whenever $f\colon X\to Y$ is an L-morphism, then $f^{-1}\colon\mathsf{Fil_{F}}(Y)\to\mathsf{Fil_{F}}(X)$ is a lattice homomorphism. This gives rise to a functor $\mathsf{Fil_{F}}\colon\mathsf{LFrm}\to\mathsf{Lat}$.
HMS duality can be restricted to lattices:
Definition 3.5.
An L-space is an HMS-space $(X,1,\curlywedge,\tau)$ such that $\{1\}$ is clopen and
$U\lor V\coloneq\{x\in X\mid\exists y\in U,z\in V\penalty 10000\ \textrm{such that}\penalty 10000\ y\curlywedge z\preceq x\},$
the filter generated by $U\cup V$, is clopen whenever $U,V$ are clopen filters.
An L-space morphism is a continuous L-morphism. We write $\mathsf{LSpace}$ for the category of L-spaces and L-space morphisms.
Theorem 3.6.
The category $\mathsf{LSpace}$ is dually equivalent to the category $\mathsf{Lat}$ of lattices.
Proof.
This is proved in [1, Theorem 2.14]. It is obtained by restricting HMS duality to lattices and join preserving morphisms. ∎
Remark 3.7.
Let us summarize the different categories and functors at play here (see also Figure 5. The categories are $\mathsf{Lat}$, the category of lattices; $\mathsf{LSpace}$, the category of L-spaces and continuous L-morphisms; and $\mathsf{LFrm}$, the category of L-frames and L-morphisms. The functors $\mathsf{Fil_{L}}\colon\mathsf{Lat}\to\mathsf{LSpace}$ and $\mathsf{ClopFil}\colon\mathsf{LSpace}\to\mathsf{Lat}$ form a duality. The functor $\mathsf{U}\colon\mathsf{LSpace}\to\mathsf{LFrm}$ is a forgetful functor, it forgets the topology, and is faithful. The functor $\mathsf{Fil_{F}}\colon\mathsf{LFrm}\to\mathsf{Lat}$ is dually adjoint to $\mathsf{U}\circ\mathsf{Fil_{L}}$.
$\mathsf{LFrm}$$\mathsf{Lat}$$\mathsf{LSpace}$$\scriptstyle{\mathsf{Fil_{F}}}$$\scriptstyle{\mathsf{Fil_{L}}}$$\scriptstyle{\mathsf{U}}$$\scriptstyle{\mathsf{ClopFil}}$
Figure 5: Summary of dualities, dual adjunctions and inclusions
Let us recall a few properties of HMS-spaces that will come in handy later.
Definition 3.8.
An ordered topological space $(X,\preceq,\tau)$ is called a Priestley space if $(X,\tau)$ is compact and whenever $x\not\preceq y$ then there is a clopen upset $U$ such that $x\in U$ and $y\notin U$.
Proposition 3.9.
Let $(X,\tau)$ be an HMS-space.
  1. (i)
    $X$ is a Priestley space.
  2. (ii)
    For any closed subset $U\subseteq X$, the set ${\uparrow}U$ is closed.
  3. (iii)
    Whenever $U$ is a closed subset of $X$ and $x\in U$, then there is a maximal point $x\preceq x$ such that $y\in U$, and a minimal point $z\preceq y$ such that $z\in U$.
  4. (iv)
    $X$ has arbitrary meets and joins.
Proof.
(i) follows directly by the definitions. (ii) and (iii) are well known properties of Priestley spaces (see [6, Lemma 11.21] and [9, Theorem 3.2.3], and hence follow from (i).
For (iv), it follows by [1, Lemma 2.8] that the intersection of principal filters is principal. Given a collection $C\subseteq X$, the generator of the principal filter
$\bigcap_{x\in C}{\uparrow}x$
is the join of $C$, denoted $\bigcurlyvee C$. ∎
We recall some more facts about this duality:
Proposition 3.10.
  1. (i)
    If $f\colon X\to Y$ is a surjective L-frame morphism, then $f^{-1}\colon\mathsf{ClopFil}(Y)\to\mathsf{ClopFil}(X)$ is an injective lattice homomorphism.
  2. (ii)
    If $h\colon K\to L$ is an injective lattice homomorphism, then $h^{-1}\colon\mathsf{Fil}(L)\to\mathsf{Fil}(K)$ is surjective.
Proof.
The first claim is trivial. For the second, assume that $h$ is injective, and let $F\in\mathsf{Fil_{L}}(K)$. Observe that $h[F]$ is closed under meets, and that ${\uparrow}h[F]$ is a filter. We show that $h^{-1}[{\uparrow}h[F]]=F$.
Clearly $F\subseteq h^{-1}[{\uparrow}h[F]]$. We show the converse. Let $a\in h^{-1}[{\uparrow}h[F]]$. Then there is $b\in F$ such that $h(b)\leq h(a)$. That is, we have $h(a\land b)=h(a)\land h(b)=h(b)$, thus $a\land b=b$ by injectivity of $h$. This immediately implies that $b\leq a$, thus $a\in F$. ∎
Using Proposition 3.10, we can translate amalgamation into a co-amalgamation property for dual spaces. To show it, we will make use of the following key separation result.
Lemma 3.11.
Let $(X,\tau)$ be an L-space. Let $U,V\subseteq X$ be two closed subsets, such that $U$ and $X\setminus V$ are filters, and $U\cap V=\emptyset$. Then there exists a clopen filter $W$ such that $U\subseteq W$ and $V\subseteq X\setminus W$.
Proof.
Fix $x\in U$. For every $y\in V$, by assumption, $x\not\preceq y$, so there is some clopen filter $W_{x,y}$ such that $x\in W_{x,y}$ and $y\not\in W_{x,y}$. Hence we have $V\subseteq\bigcup_{y\in V}X\setminus W_{x,y}$, so by compactness, there are $y_{1},\dots,y_{n},\in V$ such that $V\subseteq(X\setminus W_{x,y_{1}})\cup\dots\cup(X\setminus W_{x,y_{n}})$. Letting $W_{x}\coloneq W_{x,y_{1}}\cap\dots\cap W_{x,y_{n}}$, we have $V\subseteq X\setminus W_{x}$ and $x\in W_{x}$, and $W_{x}$ is a clopen filter.
Since $x\in W_{x}$, and this holds for each $x\in U$, we obtain $U\subseteq\bigcup_{x\in U}W_{x}$, so by compactness again, there are $x_{1},\dots,x_{n}\in U$ such that $U\subseteq W_{x_{1}}\cup\dots\cup W_{x_{n}}$. Letting $W\coloneq W_{x_{1}}\lor\dots\lor W_{x_{n}}$, we have $U\subseteq W$, $W\subseteq X\setminus V$ (because $W_{x}\subseteq X\setminus V$ and $X\setminus V$ is a filter), and $W$ is a clopen filter, as desired. ∎
3.2 Craig interpolation for weak positive logic
We will now prove superamalgamation of lattices via duality.
Theorem 3.12.
The variety of lattices has the superamalgamation property.
Proof.
Let $K,L_{1},L_{2},h_{1},h_{2}$ be a V-formation. Dually, we consider the diagram in Figure 6(b). Consider $\mathrm{Pb}(f_{1},f_{2})\coloneq\{(x,y)\mid f_{1}(x)=f_{2}(y)\}$ together with the surjective projections $\pi_{i}\colon\mathrm{Pb}(f_{1},f_{2})\to L_{i}$.
[…]
(a) Amalgam of lattices
[…]
(b) Coamalgam of L-frames
Figure 6: Amalgamation diagram and its dual
Note that $\mathrm{Pb}(f_{1},f_{2})$ is not guaranteed to be an L-space. However, it will always be an HMS-space, since products and closed substructures of HMS-spaces are still HMS-spaces. We show that $\pi_{i}$ are $L$-morphisms
First assume that $\pi_{1}(x_{1},x_{2})=1$. Then $x_{1}=1$, thus $f_{2}(x_{2})=f_{1}(x_{1})=1$, and thus $x_{2}=1$ since $f_{2}$ is an L-morphism. Now assume that $y_{1}\curlywedge z_{1}\preceq\pi_{1}(x_{1},x_{2})$. Then $y_{1}\curlywedge z_{1}\preceq x_{1}$, thus $f_{1}(x_{1})\curlywedge f_{1}(x_{1})\preceq f_{1}(x_{1})=f_{2}(x_{2})$. Since $f_{2}$ is an L-morphism, we find $y_{2}^{\prime},z_{2}^{\prime}\in Y_{2}$ such that $f_{1}(y_{1})\preceq f_{2}(y_{2}^{\prime})$, $f_{1}(z_{1})\preceq f_{2}(z_{2}^{\prime})$ and $y_{2}^{\prime}\curlywedge z_{2}^{\prime}\preceq x_{2}$. Since $f_{2}$ is onto, we also find $y_{2}^{\prime\prime},z_{2}^{\prime\prime}\in Y_{2}$ such that $f_{2}(y_{2}^{\prime\prime})=f_{1}(y_{1})$ and $f_{2}(z_{2}^{\prime\prime})=f_{1}(z_{1})$. Observe that $f_{2}(y_{2}^{\prime}\curlywedge y_{2}^{\prime\prime})=f_{1}(y_{1})$, and $f_{2}(z_{2}^{\prime}\curlywedge z_{2}^{\prime\prime})=f_{1}(z_{1})$, thus $(y_{1},y_{2}^{\prime}\curlywedge y_{2}^{\prime\prime}),(z_{1},z_{2}^{\prime}\curlywedge z_{2}^{\prime\prime})\in\mathrm{Pb}(f_{1},f_{2})$. Furthermore, we have $y_{1}\preceq\pi_{1}(y_{1},y_{2}^{\prime}\curlywedge y_{2}^{\prime\prime})$, $z_{1}\preceq\pi_{1}(z_{1},z_{2}^{\prime}\curlywedge z_{2}^{\prime\prime})$ and $(y_{1},y_{2}^{\prime}\curlywedge y_{2}^{\prime\prime})\curlywedge(z_{1},z_{2}^{\prime}\curlywedge z_{2}^{\prime\prime})\preceq(x_{1},x_{2})$. This proves that $\pi_{1}$ is an L-morphism, the proof for $\pi_{2}$ is similar.
Hence we can consider the lattice $\mathsf{Fil_{F}}(\mathrm{Pb}(f_{1},f_{2}))$, together with the injective homomorphisms $\pi_{i}^{-1}\colon L_{i}\to\mathsf{Fil_{F}}(\mathrm{Pb}(f_{1},f_{2}))$, which will be $L$-morphisms.
It is clear that the diagram commutes, so this provides an amalgam. We check that it is a superamalgam as well.
We first show the following two claims:
Claim 3.13.
Assume that $U\subseteq Y_{1}$ and $V\subseteq Y_{2}$ are two clopen filters.
  1. (i)
    ${\downarrow}f_{2}[Y_{2}\setminus V]$ is a closed subset which complement is a filter, and ${\uparrow}f_{1}[U]$ is a closed filter.
  2. (ii)
    If $\pi_{1}^{-1}[U]\subseteq\pi_{2}^{-1}[V]$, then ${\uparrow}f_{1}[U]\cap{\downarrow}f_{2}[Y_{2}\setminus V]=\emptyset$.
Proof.
The fact that ${\uparrow}f_{1}[U]$ is a filter is easy to see. To show that the complement of ${\downarrow}f_{2}[Y_{2}\setminus V]$ is a filter, take $x,y\notin{\downarrow}f_{2}[Y_{2}\setminus V]$. Suppose that $x\curlywedge y\preceq f_{2}(z)$ where $z\notin V$. Because $f_{2}$ is an L-morphism, there are $x^{\prime},y^{\prime}$ such that $x^{\prime}\curlywedge y^{\prime}\preceq z$, $x\preceq f_{2}(x^{\prime})$ and $y\preceq f_{2}(y^{\prime})$. Because $x\notin{\downarrow}f_{2}[Y_{2}\setminus V]$, we have $x^{\prime}\in V$, for otherwise, $x^{\prime}\in Y_{2}\setminus V$ so $x\preceq f_{2}(x^{\prime})$ gets us a contradiction. Similarly, $y^{\prime}\in V$ and since $V$ is a filter, $x^{\prime}\curlywedge y^{\prime}\in V$. But $V$ is a filter, so $z\in V$, a contradiction.
Both sets are closed. Indeed, $f_{2}[Y_{2}\setminus V]$ and $f_{1}[U]$ are compact, as they are images of a compact set via a continuous map, and therefore are closed. By Proposition 3.9, the sets ${\downarrow}f_{2}[Y_{2}\setminus V]$ and ${\uparrow}f_{1}[U]$ are closed.
For the second statement, assume that $y\in{\uparrow}f_{1}[U]\cap{\downarrow}f_{2}[Y_{2}\setminus V]$. Then there are points $x\in U$ and $z\in Y_{2}\setminus V$ such that $f_{1}(x)\preceq y\preceq f_{2}(z)$. Now since $f_{2}$ is surjective, there is some $t$ such that $f_{1}(x)=f_{2}(t)$. Then $f_{2}(z)\curlywedge f_{2}(t)=f_{1}(x)$ and thus $f_{2}(z\curlywedge t)=f_{1}(x)$. Hence $(x,z\curlywedge t)\in\mathrm{Pb}(f_{1},f_{2})$, and this belongs to $\pi_{1}^{-1}[U]$. By assumption, this also belongs to $\pi_{2}^{-1}[V]$, so $z\curlywedge t\in V$, and by upwards closure, $z\in V$, a contradiction. This proves the result. ∎
Now assume that $U\subseteq Y_{1}$ and $V\subseteq Y_{2}$ are clopen filters, such that $\pi_{1}^{-1}[U]\subseteq\pi_{2}^{-1}[V]$. By Claim 3.13, we obtain that there is a clopen filter $W$ such that ${\uparrow}f_{1}[U]\subseteq W$ and ${\downarrow}f_{2}[Y_{2}\setminus V]\subseteq X\setminus W$. This implies that $U\subseteq f_{1}^{-1}[W]$ and $f_{2}^{-1}[W]\subseteq V$, which proves superamalgamation. ∎
It is worth comparing this proof technique to the one exposed in [16], which is originally due to Jónsson [17]. We recall the proof technique used there.
Remark 3.14 (Jónsson’s proof of Theorem 3.12).
Let $K,L_{1},L_{2},h_{1},h_{2}$ be a V-formation. Without loss of generality, we may assume that $K$ is a sublattice of both $L_{1}$, that $L_{2}$, and $L_{1}\cap L_{2}=K$ and that $h_{1}$, $h_{2}$ are inclusions, which we denote by $\iota_{1}$, $\iota_{2}$. We have to show that there is a lattice $M$ with embeddings $p_{1}\colon L_{1}\to M$ and $p_{2}\colon L_{2}\to M$ such that $M$ is a superamalgam.
Let $\leq_{1}$ be the order on $L_{1}$ and $\leq_{2}$ be the order on $L_{2}$. We define an order $\leq$ on $L_{1}\cup L_{2}$ as the transitive closure of ${\leq_{1}}\cup{\leq_{2}}$, turning $L_{1}\cup L_{2}$ into a poset. One easily shows that $a\leq b$ iff $a,b\in L_{i}$ and $a\leq_{i}b$, or $a\in L_{i}$, $b\in L_{j}$ and there is $c\in K$ such that $a\leq_{i}c\leq_{j}b$.
A subset $F$ of $L_{1}\cup L_{2}$ is called an $(L_{1},L_{2})$-filter if it is an upset and $a,b\in F\cap L_{i}$ implies $a\land_{i}b\in F$, where $\land_{i}$ is the meet in $L_{i}$. We denote by $\mathrm{Fil}(L_{1},L_{2})$ the set of $(L_{1},L_{2})$-filters of $L_{1}\cup L_{2}$, ordered by reverse inclusion. This is a lattice, and we have embeddings $p_{i}\colon L_{i}\to\mathrm{Fil}(L_{1},L_{2})\quad a\mapsto{\uparrow}a$. One easily shows that this is a superamalgam.
The set $\mathrm{Fil}(L_{1},L_{2})$ thus defined is the order dual of the set $\mathrm{Pb}(f_{1},f_{2})$ defined in the first proof. Indeed, when the maps $h_{i}\colon K\to L_{i}$ are inclusions, the maps $f_{i}\colon Y_{i}\to X$ are are restrictions sending $F\subseteq Y_{i}$ to $F\cap K$. Therefore, the elements of $\mathrm{Pb}(f_{1},f_{2})$ are pairs of filters $(F_{1},F_{2})$ such that $F_{1}\cap K=F_{2}\cap K$. One easily verifies that if $(F_{1},F_{2})$ is such a pair, then $F_{1}\cup F_{2}$ is an $(L_{1},L_{2})$-filter; and that conversely, if $F$ is an $(L_{1},L_{2})$-filter, then $(F\cap L_{1},F\cap L_{2})$ is an element of $\mathrm{Pb}(f_{1},f_{2})$; those two constructions being each other’s inverse. As $\mathrm{Fil}(L_{1},L_{2})$ is ordered by reverse inclusion, while $\mathrm{Pb}(f_{1},f_{2})$ is ordered by pairwise inclusion, their orders are reversed.
Remark 3.15.
While Jónsson’s proof is striking by its simplicity, it should be noted that it does not straightforwardly generalize to the modal case. This can be explained by the fact that the amalgam $\mathrm{Fil}(L_{1},L_{2})$ actually belongs to the dual side. Therefore, if we were to add modalities to the initial lattices, this would correspond to adding a relation on $\mathrm{Fil}(L_{1},L_{2})$, but it is unclear how this could induce modalities on that same lattice.
On the other hand, we will see in the next section that the method presented in this paper generalizes to the modal case, with the modalities on the initial lattices corresponding to relations on the dual L-frame $\mathrm{Pb}(f_{1},f_{2})$, which in turn can be turned into modalities on $\mathsf{Fil_{F}}(\mathrm{Pb}(f_{1},f_{2}))$
4 Superamalgamation of modal lattices
4.1 Duality for modal lattices
In this section we look at superamalgamation of modal lattices, using the general strategy employed in the previous section. For that purpose, we extend the duality used in the previous section. Let us first recall the modal version of L-frames, first introduced in [1].
Definition 4.1.
A modal L-frame is a tuple $(X,1,\curlywedge,R)$ where $(X,1,\curlywedge)$ is an L-frame, and $R$ is a binary relation on $X$ such that:
  1. (i)
    if $x\preceq y$ and $y\mathrel{R}z$, there is a $w\in X$ such that $x\mathrel{R}w$ and $w\preceq z$;
  2. (ii)
    if $x\preceq y$ and $x\mathrel{R}w$, there is a $z\in X$ such that $y\mathrel{R}z$ and $w\preceq z$;
  3. (iii)
    if $(x\curlywedge y)\mathrel{R}z$, there are $u,v\in X$ such that $x\mathrel{R}u$, $y\mathrel{R}v$ and $u\curlywedge v\preceq z$;
  4. (iv)
    if $x\mathrel{R}u$ and $y\mathrel{R}v$, then $(x\curlywedge y)\mathrel{R}(u\curlywedge v)$;
  5. (v)
    for all $x\in X$, $1\mathrel{R}x$ iff $x=1$.
[…]
Figure 7: The first four conditions of a modal L-frame. Lines denote the poset order, with high nodes being bigger. Arrows denote the relation $R$.
Any modal L-frame $(X,R)$ gives rise to modal operators on the lattice $\mathsf{Fil_{F}}(X)$. Indeed, the five conditions ensure that $\Box_{R}\colon U\mapsto\{x\in X\mid R[x]\subseteq U\}$ and $\Diamond_{R}\colon U\mapsto\{x\in X\mid R[x]\cap U\neq\emptyset\}$ send filters to filters.
Definition 4.2.
A bounded L-morphism between modal L-frames $(X,R)$ and $(Y,R)$ is an L-morphism $f\colon X\to Y$ such that:
  1. (i)
    if $x\mathrel{R}y$, then $f(x)\mathrel{R}f(y)$;
  2. (ii)
    if $f(x)\mathrel{R}z$, then there is $y\in X$ such that $x\mathrel{R}y$ and $f(y)\preceq z$;
  3. (iii)
    if $f(x)\mathrel{R}z$, then there is $y\in X$ such that $x\mathrel{R}y$ and $z\preceq f(y)$.
One easily checks that if $f\colon(X,R)\to(Y,R)$ is a bounded L-morphism, then $f^{-1}\colon\mathsf{Fil_{F}}(Y)\to\mathsf{Fil_{F}}(X)$ preserves $\Box_{R}$ and $\Diamond_{R}$.
Now, to obtain a duality, a topology needs to be added to modal L-frames. This gives the following definition.
Definition 4.3.
A modal L-space is a tuple $(X,1,\curlywedge,\tau,R)$ such that:
  1. (i)
    $(X,1,\curlywedge,\tau)$ is an L-space;
  2. (ii)
    $R$ is a binary relation on $X$ such that $1\mathrel{R}x$ iff $x=1$;
  3. (iii)
    if $U$ is a clopen filter, then so are $\Box_{R}U$ and $\Diamond_{R}U$;
  4. (iv)
    for all $x,y\in X$, we have $x\mathrel{R}y$ iff for all clopen filter $U$, $x\in\Box_{R}U$ implies $y\in U$ and $y\in U$ implies $x\in\Diamond_{R}U$.
A modal L-space morphism is a continuous bounded L-morphism.
We write $\mathsf{MLSpace}$ for the category of modal L-spaces and modal L-space morphisms.
If $(X,\tau,R)$ is a modal L-space, then $(X,R)$ is a modal L-frame (see [1, Lemma 4.15]). We denote the underlying L-frame of a modal L-space by $\mathsf{U}(X)$. Let us recall some basic facts about modal L-spaces.
Theorem 4.4.
The category $\mathsf{MLSpace}$ is dually equivalent to the category $\mathsf{MLat}$ of modal lattices.
Proof.
See [1, Theorem 4.26] for a detailed proof. Here is a quick sketch.
As noted in Theorem 3.3, the functors $\mathsf{Fil_{L}}$ and $\mathsf{ClopFil}$ send lattices to L-spaces and vice versa. Given a modal lattice $(L,\Box,\Diamond)$, we define a relation $R$ on $\mathsf{Fil_{L}}(L)$ by
$x\mathrel{R}y\qquad\text{iff}\qquad\text{for all}\penalty 10000\ a\in L,\penalty 10000\ \Box a\in x\penalty 10000\ \text{implies}\penalty 10000\ a\in y\penalty 10000\ \text{and}\penalty 10000\ a\in y\penalty 10000\ \text{implies}\penalty 10000\ \Diamond a\in x.$
This turns $\mathsf{Fil_{L}}(L)$ into a modal L-space.
In turn, given a modal L-space $(X,\tau,R)$, the conditions on $R$ ensure that $\Box_{R}$ and $\Diamond_{R}$ are well defined modal operators. Moreover, one easily observes that if an L-space morphism $f\colon(X,\tau,R)\to(Y,\tau,R)$ is a bounded morphism, then its dual $f^{-1}\colon\mathsf{ClopFil}(Y)\to\mathsf{ClopFil}(X)$ preserves modalities. Similarly, if $h\colon(L,\Box,\Diamond)\to(L,\Box,\Diamond)$ is a modal lattice homomorphism, then $h^{-1}\colon\mathsf{Fil_{L}}(L)\to\mathsf{Fil_{L}}(K)$ is a bounded morphism. ∎
Proposition 4.5.
Let $(X,\tau,R)$ be an L-space. Then $R[x]$ is a closed nonempty convex set for all $x\in X$. Moreover, for any $x\in X$, the set $R[x]$ is closed under meets.
Proof.
The fact that is nonempty follows from conditions (i) and (v) of Definition 4.1.
To show that it is closed, assume that $y\not\in R[x]$. Then there is a clopen filter $F$ such that $x\in\Box_{R}F$ and $y\not\in F$, or such that $y\in F$ and $x\not\in\Diamond_{R}F$. In the former case, we have $R[x]\subseteq F$ and $y\not\in F$, so $F^{c}$ is an open neighborhood of $y$ disjoint from $R[x]$. In the later case, we have $R[x]\cap F=\emptyset$ and $y\in F$, so $F$ is an open neighborhood of $y$ disjoint from $R[x]$.
The fact that it is convex follows from that definition of a modal L-space. Indeed, assume that $y\preceq z\preceq t$ and $y,t\in R[x]$ but $z\not\in R[x]$. Then there is a clopen filter $F$ such that $x\in\Box_{R}F$ and $z\not\in F$ or such that $z\in F$ but $x\not\in\Diamond_{R}F$. In the former case, we have $R[x]\subseteq F$, thus $y\in F$, which contradicts $z\not\in F$ as $y\preceq z$. In the latter case, we have $z\in F$ and $R[x]\cap F=\emptyset$, which contradicts $t\in F$ as $t\in R[x]$ and $t\succeq z\in F$.
The last point follows from the fact that $y,z\in R[x]$ means $x\mathrel{R}y$ and $x\mathrel{R}z$, thus $x=x\curlywedge x\mathrel{R}y\curlywedge z$ by condition (iv) of Definition 4.1. ∎
Corollary 4.6.
Let $(X,\tau,R)$ be an L-space. Then for any $x,y\in X$ such that $x\mathrel{R}y$, we can find maximal and minimal elements $z,t\in R[x]$ such that $z\preceq y\preceq t$.
Proof.
This follows from Propositions 3.9 and 4.5 straightforwardly. ∎
4.2 Craig interpolation for weak positive modal logic
We are now ready to prove the main result of this paper.
Theorem 4.7.
The variety of modal lattices has the superamalgamation property.
Proof.
We proceed as in the proof of Theorem 3.12. We must define a relation on $\mathrm{Pb}(f_{1},f_{2})$, show that it gives rise to a modal L-frame, and that the projection maps $\pi_{i}$ are bounded L-morphisms. This allows us to define modalities on $\mathsf{Fil_{F}}(\mathrm{Pb}(f_{1},f_{2}))$, and shows that the maps $\pi_{i}^{-1}$ preserve those modalities.
We define the relation $R$ on $\mathrm{Pb}(f_{1},f_{2})$ as $(x_{1},x_{2})\mathrel{R}(y_{1},y_{2})$ iff $x_{1}\mathrel{R}y_{1}$ and $x_{2}\mathrel{R}y_{2}$. Let us first show that it satisfies the conditions for a modal L-frame.
  1. (i)
    Assume $(x_{1},x_{2})\preceq(y_{1},y_{2})$ and $(y_{1},y_{2})\mathrel{R}(z_{1},z_{2})$. Let $x\coloneq f_{1}(x_{1})=f_{2}(x_{2})$, $y\coloneq f_{1}(y_{1})=f_{2}(y_{2})$ and $z\coloneq f_{1}(z_{1})=f_{2}(z_{2})$. We have $x\preceq y\mathrel{R}z$, thus there is $t\in X$ such that $x\mathrel{R}t\preceq z$. By Corollary 4.6, we may assume that $t$ is minimal in $R[x]$.
    We have $f_{1}(x_{1})\mathrel{R}t$, and since $f_{1}$ is a bounded L-morphism, we can find $t^{\prime}_{1}\in Y_{1}$ such that $x_{1}\mathrel{R}t^{\prime}_{1}$ and $f_{1}(t^{\prime}_{1})\preceq t$. Because $Y_{1}$ satisfies condition (i) of a modal L-frame, we can also find $t^{\prime\prime}_{1}\in Y_{1}$ such that $x_{1}\mathrel{R}t^{\prime\prime}_{1}\preceq z_{1}$.
    Let $t_{1}\coloneq t^{\prime}_{1}\curlywedge t^{\prime\prime}_{1}$. Then $x_{1}\mathrel{R}t_{1}$ (by condition (iv) of a modal L-frame), $t_{1}\preceq z_{1}$ and $f_{1}(t_{1})\preceq t$. However, as $f_{1}$ preserves $R$, we have $f_{1}(t_{1})\in R[x]$, thus $f_{1}(t_{1})=t$ by minimality of $t$.
    Similarly, we find $t_{2}\in Y_{2}$ such that $x_{2}\mathrel{R}t_{2}$, $t_{2}\preceq z_{2}$ and $f_{2}(t_{2})=t$. Then, the pair $(t_{1},t_{2})$ is in $\mathrm{Pb}(f_{1},f_{2})$, and we have $(x_{1},x_{2})\mathrel{R}(t_{1},t_{2})\preceq(z_{1},z_{2})$.
  2. (ii)
    We proceed similarly.
  3. (iii)
    Assume $(x_{1},x_{2})\curlywedge(y_{1},y_{2})\mathrel{R}(z_{1},z_{2})$. Let $x\coloneq f_{1}(x_{1})=f_{2}(x_{2})$, $y\coloneq f_{1}(x_{1})=f_{2}(x_{2})$ and $z\coloneq f_{1}(z_{1})=f_{2}(z_{2})$. We have $x\curlywedge y\mathrel{R}z$, thus there are $u,v\in X$ such that $x\mathrel{R}u$, $y\mathrel{R}v$ and $u\curlywedge v\preceq z$. By Corollary 4.6, we may assume that $u$ is minimal in $R[x]$, and $v$ is minimal in $R[y]$.
    We have $f_{1}(x_{1})\mathrel{R}u$, and $f_{1}(y_{1})\mathrel{R}v$, and since $f_{1}$ is a bounded L-morphism, we can find $u^{\prime}_{1},v^{\prime}_{1}\in Y_{1}$ such that $x_{1}\mathrel{R}u^{\prime}_{1}$, $y_{1}\mathrel{R}v^{\prime}_{1}$, $f_{1}(u^{\prime}_{1})\preceq u$ and $f_{1}(v^{\prime}_{1})\preceq v$. Since $x_{1}\curlywedge y_{1}\mathrel{R}z_{1}$, we can also find $u^{\prime\prime}_{1},v^{\prime\prime}_{1}\in Y_{1}$ such that $x_{1}\mathrel{R}u^{\prime\prime}_{1}$, $y_{1}\mathrel{R}v^{\prime\prime}_{1}$ and $u^{\prime\prime}_{1}\curlywedge v^{\prime\prime}_{1}\preceq z_{1}$. Let $u_{1}\coloneq u^{\prime}_{1}\curlywedge u^{\prime\prime}_{1}$, and $v_{1}\coloneq v^{\prime}_{1}\curlywedge v^{\prime\prime}_{1}$. Then $x_{1}\mathrel{R}u_{1}$, $y_{1}\mathrel{R}v_{1}$, $u_{1}\curlywedge v_{1}\preceq z_{1}$, $f_{1}(u_{1})\preceq u$ and $f_{1}(v_{1})\preceq v$. From $f_{1}(u_{1})\in R[x]$, $f_{1}(u_{1})\preceq u$ and the minimality of $u$, we get $f_{1}(u_{1})=u$. Similarly, we get $f_{1}(v_{1})=v$.
    Proceeding similarly, we obtain $u_{2},v_{2}\in Y_{2}$ such that $x_{2}\mathrel{R}u_{2}$, $y_{2}\mathrel{R}v_{2}$, $u_{2}\curlywedge v_{2}\preceq z_{2}$, $f_{2}(u_{2})=u$ and $f_{2}(v_{2})=v$. Then, the pairs $(u_{1},u_{2})$ and $(v_{1},v_{2})$ are in $\mathrm{Pb}(f_{1},f_{2})$, and we have $(x_{1},x_{2})\mathrel{R}(u_{1},u_{2})$, $(y_{1},y_{2})\mathrel{R}(v_{1},v_{2})$ and $(u_{1},u_{2})\curlywedge(v_{1},v_{2})\preceq(z_{1},z_{2})$.
The last two conditions are easy to verify, since everything is defined componentwise. Now we show that the projections $\pi_{i}$ are bounded L-morphisms. We do it for $\pi_{1}$, showing that it satisfies all three conditions.
  1. (i)
    Follows straightforwardly from the definition of $R$ on $\mathrm{Pb}(f_{1},f_{2})$.
  2. (ii)
    Assume that $\pi_{1}(x_{1},x_{2})\mathrel{R}z_{1}$, that is, $x_{1}\mathrel{R}z_{1}$. Let $x\coloneq f_{1}(x_{1})=f_{2}(x_{2})$ and $z\coloneq f_{1}(z_{1})$. Then $x\mathrel{R}z$, and we can find a $t\preceq z$ that is minimal in $R[x]$. As $f_{1}(x_{1})\mathrel{R}t$, we can find $y_{1}\in Y_{1}$ and $y_{2}\in Y_{2}$ such that $x_{1}\mathrel{R}y_{1}$, $x_{1}\mathrel{R}y_{2}$, $f_{1}(y_{1})\preceq t$ and $f_{2}(y_{2})\preceq t$. But then, we have $f_{1}(y_{1}),f_{2}(y_{2})\in R[x]$, and as $t$ is minimal, $f_{1}(y_{1})=t=f_{2}(y_{2})$. Thus the pair $(y_{1},y_{2})$ is in $\mathrm{Pb}(f_{1},f_{2})$, and we have $(x_{1},x_{2})\mathrel{R}(y_{1},y_{2})$ and $\pi_{1}(y_{1},y_{2})\preceq z$.
  3. (iii)
    We proceed similarly to the previous case. ∎
Combining this with Theorem 2.17, we obtain the following.
Corollary 4.8.
Weak positive modal logic has the Craig interpolation property.
4.3 Craig interpolation in extensions
The technique used to prove the previous result generalizes to any variety $\mathsf{K}$ of modal lattices such that there is a class $\mathsf{C}$ of modal L-frames satisfying $\mathsf{U}[\mathsf{Fil_{L}}[\mathsf{K}]]\subseteq\mathsf{C}$, $\mathsf{Fil_{F}}[\mathsf{C}]\subseteq\mathsf{K}$, and such that for any co-V-formation in $\mathsf{Fil_{L}}[\mathsf{K}]$, the co-amalgam $\mathrm{Pb}$ is in $\mathsf{C}$ (recall that the functor $\mathsf{Fil_{L}}$ turns modal lattices into modal $L$-space).
In particular, if a logic $\mathcal{L}(\Gamma)$ is sound and complete with respect to a class of L-frames axiomatized by universal Horn formulas, then the above condition is satisfied for $\mathsf{K}(\Gamma)$ — proving that $\mathsf{K}(\Gamma)$ has superamalgamation and that $\mathcal{L}(\Gamma)$ has Craig interpolation.
We start by recalling from [1] the L-frame semantics for our logic.
Definition 4.9.
A modal L-model is a pair $(X,R,V)$ where $(X,R)$ is a modal L-frame and $V\colon P\to\mathsf{Fil_{F}}(X)$ is a valuation. The interpretation of formulas at a state $x$ of a modal L-model $(X,R,V)$ is defined as follows:
$(X,R,V),x$ $\Vdash\top$ $\text{always},$
$(X,R,V),x$ $\Vdash\bot$ iff $x=1,$
$(X,R,V),x$ $\Vdash p$ iff $x\in V(p),$
$(X,R,V),x$ $\Vdash\phi\land\psi$ iff $(X,V),x\Vdash\phi\penalty 10000\ \text{and}\penalty 10000\ (X,V),x\Vdash\psi,$
$(X,R,V),x$ $\Vdash\phi\lor\psi$ iff $\exists y,z\in X\penalty 10000\ \text{such that}\penalty 10000\ (X,V),y\vdash\phi\penalty 10000\ \text{and}\penalty 10000\ (X,V),z\Vdash\psi\penalty 10000\ \text{and}\penalty 10000\ y\curlywedge z\preceq x,$
$(X,R,V),x$ $\Vdash\Box\phi$ iff $\forall y\in X,\penalty 10000\ x\mathrel{R}y\penalty 10000\ \text{implies}\penalty 10000\ (X,V),y\Vdash\phi,$
$(X,R,V),x$ $\Vdash\Diamond\phi$ iff $\exists y\in X\penalty 10000\ \text{such that}\penalty 10000\ x\mathrel{R}y\penalty 10000\ \text{and}\penalty 10000\ (X,V),y\Vdash\phi.$
Observe that the conditions imposed on a modal L-frame ensure that the truth set $V(\phi)\coloneq\{x\in X\mid(X,R,V),x\Vdash\phi\}$ of any formula $\phi$ is a filter. Thus the valuation extends to a map $V\colon Fm_{P}\to\mathsf{Fil_{F}}(X)$.
Definition 4.10.
We say that a modal L-frame $(X,R)$ validates a consequence pair $\phi\trianglelefteq\psi$, written $(X,R)\Vdash\phi\trianglelefteq\psi$, if for every valuation $V$ on $X$, we have $V(\phi)\subseteq V(\psi)$.
A class $\mathsf{C}$ of L-frames validates a consequence pair $\phi\trianglelefteq\psi$, written $\Vdash_{\mathsf{C}}\phi\trianglelefteq\psi$, if for all $(X,R)\in\mathsf{C}$, we have $(X,R)\Vdash\phi\trianglelefteq\psi$.
Definition 4.11.
We say that a logic $\mathcal{L}(\Gamma)$ is sound and complete with respect to a class $\mathsf{C}$ of L-frames if for any consequence pair $\phi\trianglelefteq\psi$, we have $\Vdash_{\mathsf{C}}\phi\trianglelefteq\psi$ iff $\phi\vdash_{\Gamma}\psi$.
The following is a straightforward adaptation of the classical case:
Proposition 4.12.
Given a logic $\mathcal{L}(\Gamma)$ and a class $\mathsf{C}$ of L-frames, $\mathcal{L}(\Gamma)$ is sound and complete with respect to $\mathsf{C}$ iff $\mathsf{U}[\mathsf{Fil_{L}}[\mathsf{K}(\Gamma)]]\subseteq\mathsf{C}$ and $\mathsf{Fil_{F}}[\mathsf{C}]\subseteq\mathsf{K}(\Gamma)$.
Definition 4.13.
Given a modal L-space $(X,\tau,R)$, a clopen valuation is an assignment $V\colon P\to\mathsf{ClopFil}(X)$. A modal L-space model is a pair $(X,\tau,R,V)$ where $(X,\tau,R)$ is a modal L-space and $V$ a clopen valuation. Formulas are interpreted on a modal L-space model $(X,V)$ as on the underlying modal L-model $({\mathsf{U}(X)},V)$.
Observe that the conditions imposed on a modal L-space ensure that the truth set $V(\phi)$ of any formula $\phi$ is a clopen filter. Thus we have a valuation $V\colon Fm_{P}\to\mathsf{ClopFil}(X)$.
A modal L-space $(X,\tau,R)$ validates a consequence pair $\phi\trianglelefteq\psi$, denoted $(X,\tau,R)\Vdash\phi\trianglelefteq\psi$, if for all clopen valuation $V$, we have $V(\phi)\subseteq V(\psi)$.
Proposition 4.14.
Given a logic $\mathcal{L}(\Gamma)$, the class of modal L-spaces validating every consequence pair in $\Gamma$ is exactly $\mathsf{Fil_{L}}[\mathsf{K}(\Gamma)]$.
We will now introduce some natural extensions of weak positive modal logic.
Definition 4.15.
We define the following axioms:
$\mathsf{T}$ $\coloneq\Box p\trianglelefteq p,$ $p\trianglelefteq\Diamond p,$
$\mathsf{4}$ $\coloneq\Box p\trianglelefteq\Box\Box p,$ $\Diamond\Diamond p\trianglelefteq\Diamond p,$
$\mathsf{B}$ $\coloneq p\trianglelefteq\Box\Diamond p,$ $\Diamond\Box p\trianglelefteq p,$
$\mathsf{5}$ $\coloneq\Diamond p\trianglelefteq\Box\Diamond p,$ $\Diamond\Box p\trianglelefteq\Box p,$
$\mathsf{.2}$ $\coloneq\Diamond\Box p\trianglelefteq\Box\Diamond p.$
Definition 4.16.
These axioms correspond to first order relational properties:
itemreflexivity
$\forall x,\penalty 10000\ (x\mathrel{R}x)$,
itemtransitivity
$\forall x,y,z,\penalty 10000\ (x\mathrel{R}y\penalty 10000\ \text{and}\penalty 10000\ y\mathrel{R}z\rightarrow x\mathrel{R}z)$,
itemsymmetry
$\forall x,y,(x\mathrel{R}y\rightarrow y\mathrel{R}x)$,
itemEuclideanity
$\forall x,y,z,\penalty 10000\ (x\mathrel{R}y\penalty 10000\ \text{and}\penalty 10000\ x\mathrel{R}z\rightarrow y\mathrel{R}z)$,
itemdirectedness
$\forall x,y,z,\penalty 10000\ (x\mathrel{R}y\penalty 10000\ \text{and}\penalty 10000\ x\mathrel{R}z\rightarrow\exists t\penalty 10000\ (y\mathrel{R}t\penalty 10000\ \text{and}\penalty 10000\ z\mathrel{R}t))$.
The following theorem extends well known results for classical modal logic.
Theorem 4.17.
Let $\Gamma\subseteq\{\mathsf{T},\mathsf{4},\mathsf{B},\mathsf{5},\mathsf{.2}\}$, and let $\Theta_{\Gamma}$ be the corresponding set of first order relational properties.
  1. (i)
    The class of L-spaces axiomatized by $\Theta_{\Gamma}$ is exactly $\mathsf{Fil_{L}}[\mathsf{K}(\Gamma)]$.
  2. (ii)
    The logic $\mathcal{L}(\Gamma)$ is sound and complete with respect to the class $\mathsf{C}$ of L-frames axiomatized by $\Theta_{\Gamma}$.
Proof.
We first show that a modal L-space $(X,\tau,R)$ validating every consequence pair in $\Gamma$ also satisfies the conditions in $\Theta_{\Gamma}$. By [1, Lemma 4.33], the underlying L-frame $\mathsf{U}(X)$} validates the consequence pairs in $\Gamma$.
By the Sahlqvist algorithm described in [1, Theorem 4.39], we know that $(X,R)$ satisfies
$\forall x\penalty 10000\ \exists y_{1}\penalty 10000\ (x\mathrel{R}y_{1}\penalty 10000\ \text{and}\penalty 10000\ y_{1}\preceq x)\qquad\text{and}\qquad\forall x\penalty 10000\ \exists y_{2}\penalty 10000\ (x\mathrel{R}y_{2}\penalty 10000\ \text{and}\penalty 10000\ x\preceq y_{2})$
if $\mathsf{T}\in\Gamma$,
$\forall x,y,z\penalty 10000\ (x\mathrel{R}y\mathrel{R}z\rightarrow\exists t_{1}\penalty 10000\ (x\mathrel{R}t_{1}\penalty 10000\ \text{and}\penalty 10000\ t_{1}\preceq z))\qquad\text{and}\\ \forall x,y,z\penalty 10000\ (x\mathrel{R}y\mathrel{R}z\rightarrow\exists t_{2}\penalty 10000\ (x\mathrel{R}t_{2}\penalty 10000\ \text{and}\penalty 10000\ z\preceq t_{2})).$
if $\mathsf{4}\in\Gamma$,
$\forall x,y\penalty 10000\ (x\mathrel{R}y\rightarrow\exists z_{1}\penalty 10000\ (y\mathrel{R}z_{1}\penalty 10000\ \text{and}\penalty 10000\ x\preceq z_{1}))\qquad\text{and}\qquad\forall x,y\penalty 10000\ (x\mathrel{R}y\rightarrow\exists z_{2}\penalty 10000\ (y\mathrel{R}z_{2}\penalty 10000\ \text{and}\penalty 10000\ z_{2}\preceq x))$
if $\mathsf{B}\in\Gamma$,
$\forall x,z,y\penalty 10000\ (x\mathrel{R}z\penalty 10000\ \text{and}\penalty 10000\ x\mathrel{R}y\rightarrow\exists t_{1}\penalty 10000\ (y\mathrel{R}t_{1}\penalty 10000\ \text{and}\penalty 10000\ z\preceq t_{1}))\qquad\text{and}\\ \forall x,y,z\penalty 10000\ (x\mathrel{R}y\penalty 10000\ \text{and}\penalty 10000\ x\mathrel{R}z\rightarrow\exists t_{2}\penalty 10000\ (y\mathrel{R}t_{2}\penalty 10000\ \text{and}\penalty 10000\ t_{2}\preceq z))$
if $\mathsf{5}\in\Gamma$ (note that we swapped $y$ and $z$ in the first formula), and
$\forall x,y,z\penalty 10000\ (x\mathrel{R}y\penalty 10000\ \text{and}\penalty 10000\ x\mathrel{R}z\rightarrow\exists u,v\penalty 10000\ (y\mathrel{R}u\penalty 10000\ \text{and}\penalty 10000\ z\mathrel{R}v\penalty 10000\ \text{and}\penalty 10000\ u\preceq v))$ (1)
if $\mathsf{.2}\in\Gamma$.
For the first four, recall that in a modal L-space, the set $R[x]$ is always convex (that is, if $x\mathrel{R}t_{1},x\mathrel{R}t_{2}$, and $t_{1}\preceq y\preceq t_{2}$, then $x\mathrel{R}y$). The result follows. The axiom $\mathsf{.2}$ requires a bit more work.
Take $x,y,z\in X$ such that $x\mathrel{R}y$ and $x\mathrel{R}z$. By (1), there are $u\in R[y],v\in R[z]$ such that $u\preceq v$. Swapping $y$ and $z$ in (1), we also get $s\in R[y],t\in R[z]$ such that $t\preceq s$. We have
$u\curlywedge s\preceq v\curlywedge s\preceq v\qquad\text{and}\qquad t\curlywedge s\preceq v\curlywedge s\preceq s,$
and as $u\curlywedge v,v\in R[y]$ and $t\curlywedge s,s\in R[z]$, we get $v\curlywedge s\in R[y]\cap R[z]$ by convexity. This proves directedness.
This finishes the proof that every modal L-space in $\mathsf{Fil_{L}}[\mathsf{K}(\Gamma)]$ satisfies the conditions in $\Gamma$. It follows that $\mathsf{U}[\mathsf{Fil_{L}}[\mathsf{K}(\Gamma)]]\subseteq\mathsf{C}$. Checking that any modal L-frame in $\mathsf{C}$ validates the consequence pairs in $\Gamma$ is standard, and it follows immediately that any modal L-space satisfying the conditions in $\Theta_{\Gamma}$ validates the consequence pairs in $\Gamma$ (as $(X,\tau,R)$ validates a consequence pair if $\mathsf{U}(X)$ does). ∎
Remark 4.18.
Note that the class $\mathsf{C}$ of modal L-frames introduced in the previous theorem need not be the largest class of modal L-frame that $\mathcal{L}(\Gamma)$ is sound and complete for.
We can now prove that the varieties of modal lattices axiomatized by the axioms introduced above has the superamalgamation property.
Theorem 4.19.
Let $\Gamma\subseteq\{\mathsf{T},\mathsf{4},\mathsf{B},\mathsf{5},\mathsf{.2}\}$. Then $\mathsf{K}(\Gamma)$ has the superamalgamation property
Proof.
Let $\Theta_{\Gamma}$ be the set of first order conditions corresponding to $\Gamma$, and let $\mathsf{C}$ be the class of modal L-frames axiomatized by $\Theta_{\Gamma}$. We already know that $\mathsf{U}[\mathsf{Fil_{L}}[\mathsf{K}(\Gamma)]]\subseteq\mathsf{C}$, and that $\mathsf{Fil_{F}}[\mathsf{C}]\subseteq\mathsf{K}(\Gamma)$. It is sufficient to show that given a co-V-formation of L-spaces $X,Y_{1},Y_{2}$ and $f_{1}\colon Y_{1}\to X$, $f_{2}\colon Y_{2}\to X$, the L-frame $\mathrm{Pb}(f_{1},f_{2})$ is in $\mathsf{C}$ (that is, satisfies the conditions in $\Theta_{\Gamma}$).
If $\Gamma\subseteq\{\mathsf{T},\mathsf{4},\mathsf{B},\mathsf{5}\}$, then $\Theta_{\Gamma}$ is a set of universal Horn formulas, which implies that $\mathsf{C}$ is closed under products and substructures. As the pullback is obtained using those two constructions, we know that $\mathrm{Pb}(f_{1},f_{2})\in\mathsf{C}$.
If $\mathsf{.2}\in\Gamma$, we know that $X,Y_{1},Y_{2}$ are directed, and we show that $\mathrm{Pb}(f_{1},f_{2})$ is directed. Take $(x_{1},x_{2}),(y_{1},y_{2}),(z_{1},z_{2})\in\mathrm{Pb}(f_{1},f_{2})$ such that $(x_{1},x_{2})\mathrel{R}(y_{1},y_{2})$ and $(x_{1},x_{2})\mathrel{R}(z_{1},z_{2})$. Let $x\coloneq f_{1}(x_{1})=f_{2}(x_{2})$, $y\coloneq f_{1}(y_{1})=f_{2}(y_{2})$ and $z\coloneq f_{1}(z_{1})=f_{2}(z_{2})$. By directedness, $R[y]\cap R[z]$ is nonempty. It is a closed set, therefore we can pick a minimal point $t\in R[y]\cap R[z]$.
We have $f_{1}(y_{1})\mathrel{R}t$, thus by the bounded L-morphism condition, there is $t^{\prime}_{1}\in Y_{1}$ such that $y_{1}\mathrel{R}t^{\prime}_{1}$ and $f_{1}(t_{1})\preceq t$. As $f_{1}(z_{1})\mathrel{R}t$, we also find $t^{\prime\prime}_{1}\in Y_{1}$ such that $z_{1}\mathrel{R}t^{\prime\prime}_{1}$ and $f_{1}(t^{\prime\prime}_{1})\preceq t$. As $Y_{1}$ is directed, we find $t^{\prime\prime\prime}_{1}\in R[y_{1}]\cap R[z_{1}]$. Let $t_{1}\coloneq(t^{\prime}_{1}\curlywedge t^{\prime\prime\prime}_{1})\curlyvee(t^{\prime\prime}_{1}\curlywedge t^{\prime\prime\prime}_{1})$.
We have $f_{1}(t_{1})\preceq t$. Indeed, this follows from
$f_{1}(t^{\prime}_{1}\curlywedge t^{\prime\prime\prime}_{1})\preceq f_{1}(t^{\prime}_{1})\preceq t\qquad\text{and}\qquad f_{1}(t^{\prime\prime}_{1}\curlywedge t^{\prime\prime\prime}_{1})\preceq f_{1}(t^{\prime\prime}_{1})\preceq t.$
We have $y_{1}\mathrel{R}t_{1}$. Indeed, $t^{\prime}_{1}\curlywedge t^{\prime\prime\prime}_{1}\preceq t_{1}\preceq t^{\prime\prime\prime}_{1}$, and since $t^{\prime}_{1},t^{\prime\prime\prime}_{1}\in R[y_{1}]$, we obtain $t^{\prime}_{1}\curlywedge t^{\prime\prime\prime}_{1}\in R[y_{1}]$. Then $t_{1}\in R[y_{1}]$ by convexity. Similarly, we have $z_{1}\mathrel{R}t_{1}$.
From this, it follows that $f_{1}(t_{1})\in R[y]\cap R[z]$. As $t$ is minimal in that set, from $t\preceq f_{1}(t_{1})$ we obtain $t=f_{1}(t_{1})$. Proceeding similarly, we find $t_{2}\in Y_{2}$ such that $t_{2}\in R[y_{2}]\cap R[z_{2}]$ and $f_{2}(t_{2})=t$. The pair $(t_{1},t_{2})$ then belongs to $\mathrm{Pb}(f_{1},f_{2})$ and is such that $(y_{1},y_{2})\mathrel{R}(t_{1},t_{2})$ and $(z_{1},z_{2})\mathrel{R}(t_{1},t_{2})$, showing that $\mathrm{Pb}(f_{1},f_{2})$ is a directed L-frame. ∎
Corollary 4.20.
Let $\Gamma\subseteq\{\mathsf{T},\mathsf{4},\mathsf{B},\mathsf{5},\mathsf{.2}\}$. Then $\mathcal{L}(\Gamma)$ has the Craig interpolation property.
Remark 4.21.
One should observe that in the classical case, the directed modal logic $\mathsf{K.2}$ does not have Craig interpolation [20, Section 5.6.2]. This makes the above result quite surprising. This can be explained by the fact that when working with modal L-spaces, many quantifiers can be eliminated by picking suitable maximal or minimal elements: in short, the usual impediments to amalgamation deriving from the need for choices disappear, since these choices become canonical.
Remark 4.22.
The above results point to some general connections between classical modal logic, positive modal logic and weak positive modal logic. For instance, if $\Gamma$ is a set of positive formulas, and $\mathcal{L}(\Gamma)$, the smallest normal modal logic containing $\Gamma$, has the following properties:
  • $\mathcal{L}(\Gamma)$ has the Lyndon interpolation property, and
  • $\mathcal{L}(\Gamma)$ is conservative over its positive fragment,
then the positive fragment of $\mathcal{L}(\Gamma)$ will have the Craig interpolation property. It would be interesting to investigate this phenomenon, as well as the relationship between positive modal logics and weak positive modal logics which are their fragments.
5 Conclusions and future work
In this paper we proved that the variety of modal lattices and its “neighbours” have the superamalgamation property via a duality between modal lattices and modal L-spaces. Consequently, the corresponding systems of modal weak positive logic have the Craig interpolation property. This includes the modal weak positive logic axiomatized by the Church–Rosser axiom, whose classical analogue is known not to have the Craig interpolation property. Below we list several directions for future work.
The proofs given in this paper are closely related to proofs of Craig interpolation via bisimulation products [2, 24]. We leave it open whether similar proofs can be developed for the results presented here. Bisimulations for L-spaces have been studied in [8]. However, a general theory of bisimulations for modal L-spaces remains to be developed.
An important strengthening of the Craig interpolation property is uniform interpolation [23]. It is well known that the basic modal logic $\mathsf{K}$, as well as the intuitionistic propositional calculus $\mathsf{IPC}$, enjoy the uniform interpolation property. We leave it as an open problem whether (modal) weak positive logic has the uniform interpolation property.
As pointed out in the introduction, there are only two nontrivial varieties of lattices with the amalgamation property, and only the variety of all lattices has the superamalgamation property. We leave open the problem of obtaining a full characterization of varieties of modal lattices with the (super)amalgamation property — i.e., is it possible to obtain a Maksimova style characterization of such varieties, possibly over a natural subvariety, such as over the transitivity axiom (see Theorem 4.19)? The duality techniques developed in this paper may be crucial for achieving this goal, as Maksimova’s characterization is also based on duality theoretic methods [5, 18]. An important component of such an approach would be the use of finite meet semilattices as dual spaces in order to construct varieties that lack the superamalgamation property.
Finally, the question of amalgamation, and more generally of Craig interpolation, in modal distributive logics also deserves further investigation. Although the variety of distributive lattices lacks the superamalgamation property, we have shown in this paper that positive (distributive) logic nonetheless has the Craig interpolation property. We leave a systematic study of Craig interpolation in distributive modal logics (a.k.a. positive modal logics) as an interesting open problem.